# Weakly supervised image segmentation for defect-based grading of fresh produce

### v1

@@ 2.5 Postprocessing
In some cases, it can be ambiguous whether to treat closely located defects as a single defect or multiple defects. To address this and ensure more consistent defect separation, we introduce a simple algorithm-based postprocessing method with the following three steps: (1) Combine all defect instances of the same type into a single binary map. (2) Use a connected components algorithm (OpenCV, 2015) to generate new instances where no pixels are connected. 3) Apply a dilation operation to expand all instance masks by $d$ pixels (we choose $d=5$) in all directions. This expanded auxiliary mask is then used to identify overlapping instances, which are subsequently merged. In other words, defect instances of the same class within a distance of $2d$ pixels are merged.
Figure 3: Distribution of IoU values when comparing annotated and SAM-generated defect masks. We generally observe a high overlap, while SAM2 masks are closer aligned to the hand-annotated ones.
@@ 2.6 Deriving number and size of defects
In the context of quality estimation, we are interested in two practical indicators that can be derived from an image’s panoptic prediction: The number and size of defects. While the first can be estimated by counting the number of defect instances, an accurate metrical size estimation in a two-dimensional image is impossible due to unknown scale and distortions. Thus, we use the relative defect size, simplified as the number of visible defect pixels divided by the number of visible foreground banana pixels, as a proxy.
3 Results and Discussion
3.1 Comparison of annotated and generated masks
Our approach uses SAM to generate dense segmentation masks in a zero-shot setting (i.e., without pre-training). To assess the quality of these masks, we first compare the masks produced by different SAM models (SAM and SAM2) with hand-annotated ground truth masks (Figure 3). Overall, we observe a high overlap between the hand-annotated defect masks and those generated by both SAM models. Notably, SAM2 exhibits even greater accuracy than SAM, with 90.5% of its masks achieving an IoU score of 0.5 or higher. Based on these findings, we conclude—and later demonstrate—that our zero-shot mask generation approach is both valid and capable of producing segmentation results comparable to hand-annotated masks.
Figure 4: Hand-annotated vs. estimated defect counts per image. We predict the exact number of defects 36.5% of the time (green squares). 78.3% of the time, we predict correctly within a ±$\pm 1$ tolerance (orange squares). Generally, our model tends to predict more defects in the images compared to the annotations.
Figure 5: Annotated vs estimated defect sizes. We pairwise match those defect instances with the highest IoU agreement between annotation and estimation (minimum 0.5) and calculate their sizes relative to the corresponding foreground banana masks. Relative size prediction is generally accurate with a Pearson correlation of $r=0.96,n=829$. Each blue dot is one pair of defects, the orange line is a fitted regression line, black line is $x=y$ diagonal.
3.2 Evaluation with image segmentation metrics
@@
Next, we assess image segmentation results on validation images using the metrics outlined in Section 2.2. Table 1 presents these metrics across various experimental setups: In the first experiment, we used hand-annotated defect masks for both training and validation, achieving an average panoptic quality (PQ) of 77.5% after applying postprocessing. This serves as our baseline for assessing how segmentation quality changes when replacing hand-annotated masks with SAM-generated ones. In the second experiment, we employ SAM-generated masks for both training and validation. The results show comparable PQ scores of 77.3% (SAM) and 77.6% (SAM2). In the final experiment, SAM-generated masks are used for training, while hand-annotated masks are used for validation. Here, the PQ scores are approximately $\mathord{\sim}0.5\%$ lower for both SAM and SAM2 compared to when hand-annotated masks are used for training. In conclusion, while SAM-generated masks are not fully equivalent to hand-annotated ones—occasionally failing to capture certain edge cases—they still produce metrics that are sufficiently close to be considered a viable low-effort alternative: For our dataset, we estimated an annotation time of approx. 15 seconds per image for coarse annotations and 3-5 minutes per image for pixel-wise annotations. An additional interpretation of these results is that there is a slight but consistent difference in how our human annotator and SAM delineate defect boundaries. Nonetheless, the Maskformer model demonstrates the ability to learn from both annotation styles effectively. Lastly, it is evident that our postprocessing method is effective for improving defect masks.
3.3 Qualitative evaluation
@@
The results can be qualitatively analyzed through the example visualizations in Figure 6 showing the segmentation using a single generic defect class. In general, the masks appear sharp, particularly for bananas, and the model reliably distinguishes between fruits in the foreground or center of the image and those in the background. While the detected defects do not always align with the annotations, they are generally plausible. The model often identifies blemishes that were not annotated, likely because they were deemed too minor to be significant. This underscores the challenge of defining ground-truth defect annotations in these images.
@@ 3.4 Number and size of defects
Panoptic segmentation masks can simply be used to derive the number and size of defects. We compare the number of defects detected per image (Figure 4) and the relative defect sizes (Figure 5) by our best model (x-axes) and the human annotations (y-axes).
The model accurately matches the exact number of defects in 36.5% of cases and is off by one in 78.3% of cases. This further highlights the issue of ambiguous ground truth, as discussed earlier.
@@
For relative size analysis, we pairwise match those defect instances with the highest IoU agreement between annotation and estimation. If no agreement with at least 0.5 IoU can be established, we consider predicted defects non-matchable and exclude them from this analysis. Relative sizes are calculated by dividing the number of pixels of a specific defect instance by the aggregated number of pixels assigned to either the foreground banana class or any defect. We generally see a high agreement in predicted vs. annotated relative sizes ($r=0.96)$, while the model tends to slightly underestimate sizes, visible by the different slopes of the regression line (orange) and $x=y$ (black dashed) lines in Figure 5.
Table 1: Main results with a single defect category. All results are reported via their mean and standard deviation over the five different cross-validation splits. We compare different setups where target defect masks are either annotated (“Anno.”) or generated by SAM/SAM2. Expectedly, when Maskformer is trained on SAM-generated masks, the defect IoU is slightly lower when evaluating annotated masks. However, recall and precision are comparable to the fully annotated setting. We also demonstrate that our postprocessing (“PP”) significantly improves precision and recall of defects. Banana segmentation quality is slightly better in the fully annotated setting ($\mathord{\sim}0.81$ vs $\mathord{\sim}0.80$ mIoU). The configuration highlighted in blue is used for subsequent visualizations.
Defect masks Defects Fruits Overall Model PP train val AP AP@50 AP@75 AR IoU IoU FG IoU BG mIoU PQ Maskformer Anno. Anno. .$133\pm.022$$243\pm.027$$134\pm.027$$256\pm.036$$593\pm.032$$952\pm.005$$732\pm.034$$812\pm.010$$758\pm.010$ Maskformer Anno. Anno. .$220\pm.019$$422\pm.014$$215\pm.033$$412\pm.041$$593\pm.032$$952\pm.005$$732\pm.034$$813\pm.010$$775\pm.007$ Maskformer SAM Anno. .$128\pm.030$$234\pm.043$$128\pm.040$$238\pm.025$$546\pm.040$$951\pm.006$$731\pm.052$$800\pm.011$$752\pm.011$ Maskformer SAM Anno. .$205\pm.017$$392\pm.040$$197\pm.013$$380\pm.022$$548\pm.039$$951\pm.006$$731\pm.052$$801\pm.011$$770\pm.005$ Maskformer SAM SAM .$137\pm.030$$241\pm.045$$4.1$034 .$250\pm.027$$561\pm.038$$951\pm.006$$731\pm.052$$804\pm.012$$755\pm.010$ Maskformer SAM SAM .$220\pm.016$$404\pm.044$$2$016 .$399\pm.023$$562\pm.037$$951\pm.006$$731\pm.052$$804\pm.011$$773\pm.005$ Maskformer SAM2 Anno. .$121\pm.028$$227\pm.047$$121\pm.036$$238\pm.034$$5.5$025 .$950\pm.005$$720\pm.041$$799\pm.015$$753\pm.011$ Maskformer SAM2 Anno. .$201\pm.022$$395\pm.018$$184\pm.030$$392\pm.020$$554\pm.025$$950\pm.005$$720\pm.041$$799\pm.015$$770\pm.013$ Maskformer SAM2 SAM2 .$134\pm.039$$229\pm.050$$139\pm.054$$249\pm.044$$579\pm.028$$950\pm.005$$720\pm.041$$805\pm.015$$757\pm.012$ Maskformer SAM2 SAM2 .$225\pm.035$$405\pm.029$$221\pm.056$$415\pm.034$$579\pm.028$$950\pm.005$$720\pm.041$$805\pm.015$$776\pm.012$
3.5 Defect Categorization
We also conducted experiments using four distinct defect types instead of a single defect class (see B for details). In this setup, the model must not only detect and segment defects but also categorize them into one of four pre-defined classes. Our results clearly show that while the detection and segmentation of bruises and scars work well, a reliable categorization into one of the four predefined classes is not possible with the current approach and dataset. We hypothesize that this limitation stems from one or more of the following factors:
  1. 1.
    Ambiguous annotation: The classification of defects across the four predefined categories may be subjective to a large degree. The distinction between “old” and ”new“ defects is not always clear-cut, as the transition is gradual. Additionally, distinguishing bruises from scars can be ambiguous, especially for non-experts (see Figure 1).
  2. 2.
  3. 3.
    The four defect types are unevenly represented in our dataset (37/82/387/834). A more balanced distribution of categories and a larger number of defect samples are likely to improve categorization accuracy. We recommend collecting more examples from the two underrepresented classes for future work.
Figure 6: Example visualizations of annotated vs predicted masks using a single defect category. Left: Input Image, Mid: Annotation (SAM), Right: Maskformer Prediction. Segments are color-coded as follows: Foreground Banana, Background Banana, Defect,
4 Conclusion and Outlook
In this work, our objective was to have an efficient way to implement machine learning to count the number of defects on banana fruit and have their relative sizes. We demonstrated the use of a panoptic image segmentation model to tackle the detection and pixel-wise segmentation of bananas and surface blemishes in images. Our approach significantly reduces the effort required for hand annotation by relying on coarse labels. Specifically, we utilize a pre-trained foundation model for image segmentation (the Segment Anything Model) to generate the dense annotations necessary for training the panoptic model. This method reduces the manual annotation workload by at least a factor of 10 while maintaining high-quality results: the measured Panoptic Quality drops from 77.5% to 77.0%) w.r.t. the annotated targets.
Our results generally demonstrate reliable and plausible segmentation of both foreground and background bananas, as well as a single generic defect category. However, when compared to human annotations, the model matches the exact number of defects in only 36.5% of cases and is off by up to one defect in 78.3% of cases. While some of this discrepancy can be attributed to model performance, it is more likely due to the target annotations. Unlike for other machine learning applications where target labels are often close to a “ground truth”, the identification of significant blemishes is to a certain degree subjective. An improvement on the labeling side could be to employ multiple annotators and measure their level of agreement. The benefits of using (noisy) annotations from multiple annotators for samples that are non-trivial to label have been explored in the machine learning literature (Welinder et al., 2010; Sheng et al., 2008). Thus, we recommend that future work carries out a more sophisticated annotation process with multiple expert annotators. The disagreement between annotators could then serve as a baseline for evaluating the performance of future machine learning models.
In addition, we demonstrated that with our current method and dataset, it is not possible to reliably categorize defects into one of the four predefined classes, namely new scars, old scars, new bruises, and old bruises. A variable ground truth is likely a contributing factor, along with the skewed class distribution and the potentially low image resolution for small defect sizes. Addressing these challenges will require further investigation in future work to improve the reliability of defect categorization.
Our method demonstrates the potential for dense image segmentation in low-data scenarios and can be broadly applied to any commodity with countable, localizable defects. However, for defects that are non-countable—such as speckles, sunburn, or chemical residues—a different machine learning paradigm (such as classification) would be more applicable.
@@ Acknowledgments
The source code of all experiments, as well as pre-trained models, can be accessed on GitHub: https://github.com/manuelknott/banana-defect-segmentation.
Supplementary Materials
Appendix A Implementation details

### v2

@@ 2.5 Postprocessing
In some cases, it can be ambiguous whether to treat closely located defects as a single defect or multiple defects. To address this and ensure more consistent defect separation, we introduce a simple algorithm-based postprocessing method with the following three steps:
  1. 1.
    Combine all defect instances of the same type into a single binary map (skip if a semantic segmentation model is used).
  2. 2.
    Use a connected components algorithm [20] to generate new instances where no pixels are connected.
  3. 3.
    Apply a dilation operation to expand all instance masks by $d$ pixels (we choose $d=5$) in all directions. This expanded auxiliary mask is then used to identify overlapping instances, which are subsequently merged. In other words, defect instances of the same class within a distance of $2d$ pixels are merged.
Note that after step 1, we have regular semantic segmentation representation, and thus, using steps 2 and 3, one could also obtain panoptic results based on standard semantic segmentation outputs. However, during initial experimentation, we found that training a Maskformer [4] yields better results than a comparable semantic segmentation model (Segformer [26]). Also, we consider this postprocessing as optional when a panoptic model is used.
@@ 2.6 Deriving number and size of defects
For many real-world grading applications, one is interested in the number and size of defects, which can both be derived (or approximated) from panoptic masks. While the first can be estimated by counting the number of defect instances, an accurate metric size estimation from a single image is generally impossible due to unknown scale, perspective distortion, and lack of depth information. Thus, we use the relative defect size, simplified as the number of visible defect pixels divided by the number of visible foreground banana pixels, as a proxy.
3 Results
3.1 SAM defect masks
Figure 2: Agreement between human-annotated and SAM-predicted masks by defect size (number of pixels). The x-axis shows binned size categories for defect sizes in pixels (per annotated mask), while $n$ denotes the number of samples in each bin. The y-axis shows the agreement between annotated and predicted masks (IoU). SAM fails to align with human annotations for small ($<\!100$ pixels) and very small ($<\!10$ pixels) defects. To understand the failure cases of larger defects ($>\!100$ pixels), we added exemplary visualizations in Figure 3. Those cases often show thin, long scars. We can also see that a larger SAM model size (ViT-H) leads to consistently higher agreement with human annotations across all defect sizes. An additional analysis of the impact of SAM model sizes can be found in Figure S.2.
Figure 3: SAM failure cases. Exemplary examples where SAM (ViT-L) fails to align with the annotated masks. We picked examples with the lowest IoU where the annotated mask size is $>\!100$ pixels. The examples shown illustrate that long, thin scars are difficult to capture. Only one defect per image is shown.
How well do SAM-predicted masks align with human annotations? We use SAM in a zero-shot manner without any fine-tuning. Thus, it is not self-evident that all defects with different shapes and sizes align closely with human annotations. Also, the choice of the SAM backbone might have an influence on predictions as model capacity usually correlates with model accuracy. Figure 3 compares the agreement between hand-annotated and SAM-generated masks for two different SAM backbone sizes (ViT-L and ViT-H) and for different defect sizes defined by the number of pixels in the annotated mask. First, we can see that a larger SAM model size (ViT-H) leads to consistently higher agreement with human annotations across all defect sizes. This finding is confirmed by an additional analysis where we plot the IoU distributions for three different SAM backbones (Figure S.2). However, we see that the alignment is generally high, and failure cases are long-tail outliers for all model backbones. A second observation from Figure 3 is that SAM fails to align with human annotations for small ($<\!100$ pixels) and very small ($<\!10$ pixels) defects. To understand the failure cases of larger defects ($>\!100$ pixels), we added exemplary visualizations in Figure 3. In those cases, which often show long, thin scars, SAM produces masks that are too large.
3.2 Panoptic predictions
Table 1: Image segmentation metrics. All results are reported via their mean and standard deviation over the five different cross-validation splits. We compare different setups where target defect masks are either annotated (“Anno.”) or generated by SAM (-L/-H refers to the size of the ViT backbone). Expectedly, when Maskformer is trained on SAM-generated masks, the defect IoU is slightly lower when evaluating annotated masks. In the same comparison, recall and precision are degrading less strongly. Postprocessing (“PP”) significantly improves precision and recall of defects. Using SAM-generated instead of annotated training masks, leads to identical PQ ($\mathord{\sim}0.779$), while mIoU is slightly worse ($\mathord{\sim}0.808$ vs $\mathord{\sim}0.810$). A larger SAM backbone size does not lead to an improvement in downstream segmentation quality. The configuration highlighted in blue is used for subsequent visualizations.
Defect masks Defects Fruits Overall Model PP train val AP AP@50 AP@75 AR IoU IoU FG IoU BG mIoU PQ Maskformer Anno. Anno. .$153\pm.040$$263\pm.066$$163\pm.050$$262\pm.051$$563\pm.037$$956\pm.006$$750\pm.050$$810\pm.018$$764\pm.010$ Maskformer Anno. Anno. .$223\pm.030$$413\pm.059$$2$040 .$397\pm.034$$563\pm.037$$956\pm.006$$750\pm.050$$810\pm.018$$779\pm.010$ Maskformer SAM-L Anno. .$129\pm.011$$231\pm.019$$134\pm.017$$243\pm.024$$559\pm.024$$952\pm.007$$746\pm.056$$808\pm.018$$760\pm.008$ Maskformer SAM-L Anno. .$217\pm.020$$405\pm.038$$211\pm.024$$393\pm.035$$562\pm.028$$952\pm.007$$746\pm.056$$808\pm.018$$779\pm.008$ Maskformer SAM-L SAM-L .$137\pm.015$$232\pm.015$$147\pm.023$$253\pm.025$$574\pm.022$$952\pm.007$$746\pm.056$$811\pm.018$$762\pm.006$ Maskformer SAM-L SAM-L .$231\pm.023$$409\pm.050$$235\pm.026$$410\pm.039$$576\pm.025$$952\pm.007$$746\pm.056$$812\pm.018$$782\pm.007$ Maskformer SAM-H Anno. .$150\pm.016$$257\pm.027$$160\pm.021$$251\pm.033$$547\pm.063$$952\pm.008$$733\pm.045$$801\pm.023$$759\pm.007$ Maskformer SAM-H Anno. .222±.028plus-or-minus.222.028.$2$028 .$418\pm.056$$211\pm.031$$384\pm.048$$548\pm.064$$952\pm.008$$733\pm.045$$801\pm.023$$7$010 Maskformer SAM-H SAM-H .$158\pm.022$$255\pm.024$$175\pm.031$$258\pm.033$$573\pm.067$$952\pm.008$$733\pm.045$$807\pm.024$$763\pm.007$ Maskformer SAM-H SAM-H .$242\pm.031$$422\pm.048$$251\pm.043$$403\pm.058$$574\pm.068$$952\pm.008$$733\pm.045$$808\pm.024$$783\pm.011$
@@ 3.2 Panoptic predictions
Next, we assess image segmentation results on validation images using the metrics outlined in Section 2.4. Table 1 presents these metrics across various experimental setups: In the first experiment, we used hand-annotated defect masks for both training and validation, achieving an average panoptic quality (PQ) of 77.9% after applying postprocessing. This serves as our baseline for assessing how segmentation quality changes when replacing hand-annotated masks with SAM-generated ones. In the second experiment, we employ SAM-generated masks for both training and validation. The results show a slightly higher PQ score of 78.2% (ViT-L backbone). In the final experiment, SAM-generated masks are used for training, while hand-annotated masks are used for validation. One can see a slight drop in detection metrics (AP: $-0.6\%$, AR: $-0.4\%$, IoU: $-0.1\%$), while the overall PQ score is almost identical when SAM-generated instead of hand-annotated masks are used for training.
While we previously showed that the larger SAM variant (ViT-H) has a higher defect mask agreement with hand annotations (see Figure 3), this does not, somewhat counterintuitively, lead to better segmentation metrics. This suggests that the downstream model (MaskFormer, in our case) can compensate for a certain degree of inaccuracy in the training masks.
Across all settings, IoU scores for segmenting foreground bananas are consistently high ($>95\%$), while those for background bananas are noticeably lower ($\mathord{\sim}74\%$). This suggests that the model rarely confuses the two classes—if it did, both scores would be low. Instead, we conclude that background bananas are occasionally missed in the annotations. Qualitative comparisons support this interpretation (see, for example, Figure 6, second and fourth row).
Furthermore, the postprocessing described in Section 2.5 yields a significant improvement in AP across all settings.
@@
The panoptic results can be qualitatively analyzed through the example visualizations in Figure 6 showing the segmentation using a single generic defect class. In general, the masks appear sharp, particularly for bananas, and the model reliably distinguishes between fruits in the foreground or center of the image and those in the background. While the detected defects do not always align with the annotations, they are generally plausible. The model often identifies blemishes that were not annotated, likely because they were deemed too minor to be significant. This underscores the challenge of defining ground-truth defect annotations in these images.
Figure 4: Annotated versus predicted defect counts per image. We predict the exact number of defects 36.2% of the time (green squares). 76.2% of the time, we predict correctly within a ±$\pm 1$ tolerance (orange squares). Generally, our model tends to predict more defects in the images compared to the annotations.
Figure 5: Annotated versus predicted defect sizes. We pairwise match those defect instances with the highest IoU agreement between annotation and estimation (minimum 0.5) and calculate their sizes relative to the corresponding foreground banana masks (non-matchable defects are excluded in this analysis). Relative size prediction is generally accurate with a Pearson correlation of $r=0.96,n=793$. Each blue dot is one pair of defects, the orange line is a fitted regression line, and the black dashed line is the $x=y$ diagonal.
Lastly, we also conducted experiments using four distinct defect types instead of a single defect class. We found that defect categorization is very unreliable and attribute this mainly to ambiguous annotations. Thus, we do not recommend using our dataset for defect categorization at this point (see Appendix D for a more detailed discussion).
Figure 6: Example visualizations of annotated vs predicted masks. Left: Input Image, Mid: Annotation, Right: Maskformer Prediction. Red rectangles enclose defect instances. Segments are color-coded as follows: Foreground Banana, Background Banana, Defect,
@@ 3.3 Number and size of defects
Panoptic segmentation masks can simply be used to derive the number and size of defects. We compare the number of defects detected per image (Figure 4) and the relative defect sizes (Figure 5) by our best model (x-axes) and the human annotations (y-axes). The model matches the exact number of defects in 36.2% of cases and is off by one in 76.2% of cases. For relative size analysis, we pairwise match those defect instances with the highest IoU agreement between annotation and estimation. If no agreement with at least 0.5 IoU can be established, we consider predicted defects non-matchable and exclude them from this analysis. Relative sizes are calculated by dividing the number of pixels of a specific defect instance by the aggregated number of pixels assigned to either the foreground banana class or any defect. We generally see a high agreement in predicted vs. annotated relative sizes ($r=0.96)$, while the model tends to slightly underestimate sizes, visible by the different slopes of the regression line (orange) and $x=y$ (black dashed) lines in Figure 5.
4 Discussion and Limitations
In this work, we lay the foundation for the automated, defect-based grading of fresh produce using limited samples and annotations. We present a carefully curated dataset comprising 476 real-world images of banana bunches and 1,440 annotated surface defects covering a wide range of sizes, shapes, and categories.
Further, we highlight the trade-off when using SAM for generating dense image masks from coarse annotations. Specifically, we estimate that the manual annotation workload can be reduced by at least a factor of 10 while image segmentation metrics for the downstream panoptic segmentation task experience only a moderate decrease (Defect IoU 56.3%→56.2%), Defect AP 22.3%→21.7%, Defect AR 39.7%→39.3%). We train a panoptic segmentation model and propose both a postprocessing algorithm and a procedure to extract grading-relevant metrics—such as defect count and relative size—from the model’s predictions. While it can be questioned if the current model is sufficient for practical deployment, it shows promise. Compared to human annotations, the model predicts the exact number of defects in 36.2% of cases and is off by at most one defect in 76.2% of cases. Although part of this discrepancy can be attributed to model limitations, we identify several avenues for improvement for future work, which we outline in Section 5.
Lastly, it is important to note that countable and sizable defects represent only a subset of relevant grading criteria. For bananas, other defect types—such as speckles, sunburn, chemical residues, or malformed shapes—may require fundamentally different approaches, such as classification or ordinal regression.
5 Future Work
Improving model performance
We see several ways to improve model performance on our dataset:
  • Model choice: The landscape of deep learning model architectures evolves rapidly. Our approach will likely benefit from better SAM variants and improved panoptic segmentation architectures. SAM 2 [21] and Mask2Former [5] are the logical next candidates to be evaluated in our framework.
  • Fine-tune SAM: Instead of using SAM in a zero-shot manner, one could fine-tune the model on domain data to achieve better segmentation masks for domain-specific instances. This was already successfully done in different contexts [see e.g., 11]. We hypothesize that this will help with the issues of thin, long scars that can be seen in Figure 3.
  • Categorized segmentation with SAM: Instead of our current two-model setup, adapting SAM to directly support categorized semantic or panoptic segmentation in a zero- or few-shot manner would provide a more elegant solution. Future research will likely explore such extensions.
  • As we have highlighted in Figure 3, SAM fails with very small defect sizes. We expect that small defect predictions will improve when increasing the used image resolution (we use 10242).
Improving data set creation
We also want to highlight the issue of ambiguity in expert annotations for defect segmentation and, in particular, categorization. We would like to encourage a broader discussion on this challenge. One potential improvement would be to involve multiple annotators and quantify their agreement. The value of using (noisy) annotations from multiple annotators, especially for non-trivial labeling tasks, has been well documented in the machine learning literature [25, 23]. We, therefore, recommend that future work adopt more sophisticated annotation protocols involving multiple experts, using inter-annotator disagreement as a baseline for evaluating machine learning models.
@@ Acknowledgments
Appendix A Dataset visualizations
Figure S.1: Defect examples. Example images of banana surface defects we aim to detect in this study. Bruises usually result from a dull impact, while scars are caused by a sharp impact. Old defects are usually darker than new defects as they oxidize over time.
Appendix B Implementation details